OpenStudy (anonymous):

What is the range of the exponential -(1/3)^x-2 ?

OpenStudy (unklerhaukus):

Which do you mean?\[f(x) = -(1/3)^{x-2}\]or\[f(x) = -(1/3)^x-2\]

OpenStudy (anonymous):

the first one

OpenStudy (unklerhaukus):

when x is negative and large in magnitude say negative infinity \[f(-\infty) \to -(1/3)^{-\infty-2}\to -(1/3)^{-\infty}\to-(3)^\infty \to -\infty\]

OpenStudy (unklerhaukus):

when x is large and positive \[f(\infty)\to-(1/3)^{\infty-2}\to-(1/3)^\infty\to\]

OpenStudy (unklerhaukus):

what does the function trend toward for a large positive input?

OpenStudy (anonymous):

0

OpenStudy (unklerhaukus):

good, also \[f(0)=-(1/3)^{0-2}=-(1/3)^0=-1\] so the function gives values between negative infinity and zero